The force(f),velocity(v) and time(t) are takenas fundamental units then the dimensions of mass are?

Grade:

1 Answers

Hasan Naqvi
9 years ago
[M] = [MLT-2]a [LT-1]b [T]c
[M] = [Ma La+b T-2a-b+c]
a=1
 
a+b=0
b=-a=-1
 
-2a-b+c=0 
-2-(-1)+c=0
-2+1+c=0
c=1
 
therefore, mass= f*t/v